ROTATIONAL SPECTRUM OF CH_{2}FI FROM 5 GHZ UP TO 1 THZ: ACCURATE SPECTROSCOPIC AND HYPERFINE PARAMETERS

\underline{CRISTINA PUZZARINI}, GABRIELE CAZZOLI, Dipartimento di Chimica ”G. Ciamician”, Università di Bologna, I-40126 Bologna, Italy; JUAN CARLOS LÓPEZ, JOSÉ LUIS ALONSO, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005, Valladolid, Spain; AGOSTINO BALDACCI, ALESSANDRO BALDAN, Dipartimento di Chimica Fisica, Università “Ca’ Foscari” Venezia, D.D. 2137, I-30123 Venezia, Italy; STELLA STOPKOWICZ, LAN CHENG, JÜRGEN GAUSS, Institut für Physikalische Chemie, Universität Mainz, D-55099 Mainz, Germany.

Guided by theoretical predictions, the rotational spectrum of fluoroiodomethane, CH_{2}FI, has been been recorded and assigned. Three different spectrometers have been employed, a Fourier Transform Microwave spectrometer, a Millimeter- /Submillimeter- wave Spectrometer, and a THz spectrometer, thus allowing to record a huge portion of the rotational spectrum, from 5 GHz up to 1 THz, and to accurately determine the ground-state rotational and centrifugal-distortion constants. The hyperfine structure of the rotational spectrum has been investigated by means of the Fourier Transform Microwave spectrometer and the Lamb-dip technique in the millimeter-/submillimeter-wave region, thus allowing the determination of the complete iodine quadrupole-coupling tensor and of the diagonal elements of the iodine spin-rotation tensor. Regarding the quantum-chemical calculations, inclusion of relativistic effects turned out to be essential for obtaining reliable and quantitative predictions for experiment, and they have been accounted for either by means of second-order direct perturbation theory or via a spin-free approach based on the Dirac Coulomb Hamiltonian, both in combination with coupled-cluster techniques to treat electron correlation and sufficiently large basis sets.